OpenStudy (anonymous):

One number is 12 greater than another.The product of the numbers is 28.Find the numbers. There are two pairs of numbers

OpenStudy (anonymous):

a - b = 12 a * b = 28

OpenStudy (anonymous):

figure it out

OpenStudy (lilg132):

x = 12 + y x * y = 28 use the method of substitution do you know how to do that?

OpenStudy (anonymous):

one pair both numbers are positive the other both number are negitive

OpenStudy (lilg132):

yes use the method of substitution

OpenStudy (anonymous):

OMG just somebody give me the fluttering answer please

OpenStudy (anonymous):

OMG no, we're here to teach not give answers out